# TacMan-Turbo: Proactive Tactile Control for Robust and Efficient Articulated Object Manipulation

### v2

@@ 6.1 Comparison with Tac-Man
Although both approaches share a similar state representation, they address the manipulation problem fundamentally differently. Tac-Man adopts a reactive strategy, prioritizing compensation for the current contact deviation—denoted as the deviation at step $i$. This approach temporarily halts object manipulation and focuses exclusively on deviation correction. Consequently, the process naturally divides into two distinct stages: execution and recovery, as evident in Fig. 6(a) and the zig-zag trajectories observed in Fig. 11. While this formulation enhances robustness by ensuring timely compensation, it introduces inefficiencies, as the recovery stage does not directly advance the manipulation task. Additionally, transitions between execution and recovery stages produce non-smooth trajectories, as highlighted in Fig. 7(a) and Fig. 11. Although Tac-Man incorporates multi-step exploration in real deployments to mitigate these limitations, the fundamental drawbacks of the reactive formulation remain.
In contrast, TacMan-Turbo’s core motivation explores whether robust manipulation can be achieved without interrupting the object interaction process. This question drives the development of TacMan-Turbo’s proactive formulation. Unlike reactive approaches, this formulation must account not only for the current $i$-step contact deviation but also anticipate the impact of the future $i+1$-step contact deviation caused by continuous, non-halting manipulation, as described in Eq. 5.
This noncausal formulation, which incorporates future state information, introduces challenges that previous approaches—where contact deviation served solely as an error signal—cannot address. To overcome this, TacMan-Turbo leverages the local kinematics of the object inherently embedded within contact deviations. By extracting and utilizing this structural information, TacMan-Turbo enables reasonable prediction of future states, effectively solving the noncausal problem.
This proactive formulation streamlines the previous execution-recovery cycle into a seamless and consistent manipulation process. As a result, it delivers significant performance gains, including an order-of-magnitude improvement in time efficiency, consistently high action efficiency, and smooth trajectories free from acceleration-deceleration cycles—all while maintaining manipulation effectiveness. Furthermore, this framework maintains robustness against unexpected disturbances that frequently occur in human-centric environments while continuing the manipulation process without stopping for reactive compensation, as shown in Fig. 13.
@@
Figure 14: Effects of base velocity direction. We investigate how initial directional alignment impacts manipulation performance. (a) The simulation setup measures the angle $\theta$ between the correct interaction direction (green arrow) and the initial motion direction $\boldsymbol{u}_{0}^{g_{i}}$ (orange arrow). (b) Time efficiency results reveal significant performance deterioration as $\theta$ increases in magnitude, with optimal performance near $\theta=0^{\circ}$. (c) Action efficiency peaks when initial direction closely aligns with the correct interaction direction and gradually decreases with increasing angular deviation. (d) The temporal progression further illustrates how different initial angles affect performance throughout the manipulation process, with properly aligned initial directions maintaining consistently higher efficiency. These findings highlight the critical importance of directional alignment in manipulation tasks.
6.2 Base Velocity Effects and Acquisition
The performance of TacMan-Turbo depends critically on the initial base velocity $\boldsymbol{u}_{0}^{g_{i}}$ introduced in Section 3.2. While our primary experiments utilized manually assigned values for this parameter, understanding both its effects on system performance and methods for its automated acquisition has significant implications for practical deployment. We analyze the magnitude and direction components separately to provide a more structured understanding of their respective influences.
@@ 6.2 Base Velocity Effects and Acquisition
Our implementation employed vision-based tactile sensors limited by their image sensors to approximately $100~\mathrm{Hz}$ sampling frequency. Alternative sensing technologies, such as capacitive or 6D force/torque sensors, can operate at frequencies exceeding $1000~\mathrm{Hz}$ [1], suggesting promising opportunities for substantial efficiency gains through TacMan-Turbo adaptation to these advanced modalities.
@@
The direction of base velocity profoundly influences manipulation efficiency. Greater alignment between $\boldsymbol{u}_{0}^{g_{i}}$ and the object’s optimal interaction direction yields higher first-step action efficiency (as defined in Section 4.1). This initial efficiency advantage propagates through subsequent steps, potentially maintaining elevated performance throughout the entire manipulation process.
To quantify this relationship, we conducted systematic simulations across 50 prismatic-joint objects (Fig. 14(a)), varying the angle $\theta$ between $\boldsymbol{u}_{0}^{g_{i}}$ and the optimal direction. The results confirm our theoretical expectations: both time efficiency (Fig. 14(b)) and action efficiency (Fig. 14(c)) reach maximum values at $\theta=0$ and deteriorate as $\theta$ increases. Furthermore, Fig. 14(d) reveals that per-step action efficiency maintains remarkable consistency throughout the manipulation process, closely tracking the first-step efficiency and achieving its peak at perfect alignment ($\theta=0$).
These findings not only validate our theoretical framework but also highlight the critical importance of appropriate base velocity direction selection. They further suggest promising research directions for automating preliminary directional information acquisition. Recent advances in vision-based direction estimation [31] offer particularly promising pathways. While such methods inevitably introduce some estimation noise, they enhance TacMan-Turbo’s operational autonomy. Moreover, our framework’s inherent robustness could effectively compensate for moderate directional inaccuracies, as demonstrated by the gradual performance degradation shown in Fig. 14(b).
7 Conclusion and Future Work
This paper introduces TacMan-Turbo, a novel proactive control framework that successfully resolves the persistent trade-off between efficiency and effectiveness in articulated object manipulation. By fundamentally reframing contact deviations as valuable kinematic information rather than mere error signals, our approach enables proactive control adjustments that significantly enhance manipulation efficiency while maintaining robust performance across diverse articulated mechanisms.
Our comprehensive evaluation—encompassing 200 simulated articulated objects with varied kinematic structures and multiple real-world experiments across fundamental joint types—demonstrates that TacMan-Turbo achieves perfect manipulation success rates (100%) while delivering substantial improvements in time efficiency, action efficiency, and trajectory smoothness compared to previous approaches. The statistical significance of these performance gains (p-values <0.0001<0.0001) confirms that TacMan-Turbo effectively addresses the effectiveness-efficiency trade-off that has challenged existing manipulation methods.
@@ 7 Conclusion and Future Work
Several promising directions for future research emerge from this work. First, developing adaptive methods for base velocity optimization could further enhance efficiency by dynamically adjusting control parameters in response to object characteristics. Second, extending the framework to support diverse tactile sensing modalities would broaden its applicability across robot platforms with different sensing capabilities. Finally, incorporating learning-based approaches for predicting object behavior could enable even more anticipatory control strategies. These enhancements will not only further boost the framework’s efficiency but also facilitate its application in increasingly complex, unstructured real-world environments.
@@ Acknowledgment
Our sincere thanks go to Yuanhong Zeng (UCLA) and Qian Long (UCLA) for their suggestion in driving Kinova, Lei Yan (LeapZenith AI Research) for mechanical design support, Yida Niu (PKU) for demo shooting, and Ms. Hailu Yang (PKU) for her assistance in procuring every piece of raw material necessary for this research. This work is supported in part by the National Science and Technology Major Project (2022ZD0114900), the National Natural Science Foundation of China (62376009), the Beijing Nova Program, the State Key Lab of General AI at Peking University, the PKU-BingJi Joint Laboratory for Artificial Intelligence, and the National Comprehensive Experimental Base for Governance of Intelligent Society, Wuhan East Lake High-Tech Development Zone.

### v3

@@ VI-A Comparison with Tac-Man
Although both approaches share a similar state representation, they address the manipulation problem fundamentally differently. Tac-Man adopts a reactive strategy, prioritizing compensation for the current contact deviation—denoted as the deviation at step $i$. This approach temporarily halts object manipulation and focuses exclusively on deviation correction. Consequently, the process naturally divides into two distinct stages: execution and recovery, as evident in Fig.˜6(a) and the zig-zag trajectories observed in Fig.˜11. While this formulation enhances robustness by ensuring timely compensation, it introduces inefficiencies, as the recovery stage does not directly advance the manipulation task. Additionally, transitions between execution and recovery stages produce non-smooth trajectories, as highlighted in Fig.˜7(a) and Fig.˜11. Although Tac-Man incorporates multi-step exploration in real deployments to mitigate these limitations, the fundamental drawbacks of the reactive formulation remain.
In contrast, TacMan-Turbo’s core motivation explores whether robust manipulation can be achieved without interrupting the object interaction process. This question drives the development of TacMan-Turbo’s proactive formulation. Unlike reactive approaches, this formulation must account not only for the current $i$-step contact deviation but also anticipate the impact of the future $i+1$-step contact deviation caused by continuous, non-halting manipulation, as described in Eq.˜4.
This non-causal formulation requires estimating future object states—a challenge that reactive approaches cannot address, as they treat contact deviations solely as instantaneous spatial error signals. TacMan-Turbo overcomes this through a temporal perspective: by analyzing sequential contact observations as time-series data, we extract object local kinematic information inherently embedded within tactile signals. This temporal interpretation enables TacMan-Turbo to predict future object states without requiring pre-specified kinematic models, allowing proactive velocity modulation that minimizes predicted future deviations.
This proactive formulation streamlines the previous execute-and-recover cycle into a continuous control loop. As a result, it delivers significant performance gains, including an order-of-magnitude improvement in time efficiency, consistently high action efficiency, and smooth trajectories free from acceleration-deceleration cycles—all while maintaining manipulation effectiveness. Furthermore, this framework maintains robustness against unexpected disturbances that frequently occur in human-centric environments while continuing the manipulation process without stopping for reactive compensation, as shown in Fig.˜13 and Supplementary Video S5.
@@
Figure 14: Effects of base velocity direction. We investigate how initial directional alignment impacts manipulation performance. (a) The simulation setup measures the angle $\theta$ between the correct interaction direction (green arrow) and the initial motion direction $\boldsymbol{u}_{0}^{g_{i}}$ (orange arrow). (b) Time efficiency results reveal significant performance deterioration as $\theta$ increases in magnitude, with optimal performance near $\theta=0^{\circ}$. (c) Action efficiency peaks when the initial direction closely aligns with the correct interaction direction and gradually decreases with increasing angular deviation. (d) The temporal progression further illustrates how different initial angles affect performance throughout the manipulation process, with properly aligned initial directions maintaining consistently higher efficiency. These findings highlight the critical importance of directional alignment in manipulation tasks. Box plots follow conventions established in Fig.˜4.
VI-B Base Velocity Effects and Acquisition
The performance of TacMan-Turbo depends critically on the initial base velocity $\boldsymbol{u}_{0}^{g_{i}}$ introduced in Sec.˜III-B. While our primary experiments utilized manually assigned values for this parameter, understanding both its effects on system performance and methods for its automated acquisition has significant implications for practical deployment. We analyze the magnitude and direction components separately to provide a more structured understanding of their respective influences.
@@ VI-B Base Velocity Effects and Acquisition
Our implementation employed vision-based tactile sensors limited by their image sensors to approximately $120\text{\,}\mathrm{Hz}$ sampling frequency. Alternative sensing technologies, such as capacitive or 6D force/torque sensors, can operate at frequencies exceeding $1000\text{\,}\mathrm{Hz}$ [92], suggesting promising opportunities for substantial efficiency gains through TacMan-Turbo adaptation to these advanced modalities.
@@
The direction of base velocity profoundly influences manipulation efficiency. Greater alignment between $\boldsymbol{u}_{0}^{g_{i}}$ and the object’s optimal interaction direction yields higher first-step action efficiency (as defined in Sec.˜IV-A). This initial efficiency advantage propagates through subsequent steps, potentially maintaining elevated performance throughout the entire manipulation process.
To quantify this relationship, we conducted systematic simulations across 50 prismatic-joint objects (Fig.˜14(a)), varying the angle $\theta$ between $\boldsymbol{u}_{0}^{g_{i}}$ and the optimal direction. The results confirm our theoretical expectations: both time efficiency (Fig.˜14(b)) and action efficiency (Fig.˜14(c)) reach maximum values at $\theta=0$ and deteriorate as $\theta$ increases. Furthermore, Fig.˜14(d) reveals that per-step action efficiency maintains remarkable consistency throughout the manipulation process, closely tracking the first-step efficiency and achieving its peak at perfect alignment ($\theta=0$).
These findings not only validate our theoretical framework but also highlight the critical importance of appropriate base velocity direction selection. They further suggest promising research directions for automating preliminary directional information acquisition. Recent advances in vision-based direction estimation [19, 20, 21] offer particularly promising pathways. While such methods inevitably introduce some estimation noise, they enhance TacMan-Turbo’s operational autonomy. Moreover, our framework’s inherent robustness could effectively compensate for moderate directional inaccuracies, as demonstrated by the gradual performance degradation shown in Fig.˜14(b).
VII Conclusion and Future Work
This paper introduces TacMan-Turbo, a novel proactive control framework that successfully mitigates the persistent trade-off between efficiency and effectiveness in articulated object manipulation. By fundamentally reframing contact deviations as valuable kinematic information rather than mere error signals, our approach enables proactive control adjustments that significantly enhance manipulation efficiency while maintaining robust performance across diverse articulated mechanisms.
[…]
@@ VII Conclusion and Future Work
Several promising directions for future research emerge from this work. First, developing adaptive methods for base velocity optimization could further enhance efficiency by dynamically adjusting control parameters in response to object characteristics. Second, extending the framework to support diverse tactile sensing modalities would broaden its applicability across robot platforms with different sensing capabilities. Finally, the current research primarily focuses on rigid, single-joint articulated objects. Third, while our current focus is on rigid, single-joint articulated objects, effectively handling soft, multi-part articulated systems in cluttered environments remains an important and promising challenge. These enhancements will not only further boost the framework’s efficiency but also facilitate its application in increasingly complex, unstructured real-world environments.
@@ Acknowledgment
Our sincere thanks go to Yuanhong Zeng (UCLA) and Qian Long (UCLA) for their suggestion in driving Kinova, Lei Yan (LeapZenith AI Research) for mechanical design support, Yida Niu (PKU) for demo shooting, Saiyao Zhang (UCAS) for simulation support, and Ms. Hailu Yang (PKU) for her assistance in procuring every piece of raw material necessary for this research.